# Uni-Hema: Unified Model for Digital Hematopathology

### v1

@@ Abstract
Digital hematopathology requires cell-level analysis across diverse disease categories, including malignant disorders (e.g., leukemia), infectious conditions (e.g., malaria), and non-malignant red blood cell disorders (e.g., sickle cell disease). Whether single-task, vision-language, WSI-optimized, or single-cell hematology models, these approaches share a key limitation: they cannot provide unified, multi-task, multi-modal reasoning across the complexities of digital hematopathology. To overcome these limitations, we propose Uni-Hema, a multi-task, unified model for digital hematopathology integrating detection, classification, segmentation, morphology prediction, and reasoning across multiple diseases. Uni-Hema leverages 46 publicly available datasets, encompassing over 700K images and 21K question-answer pairs, and is built upon Hema-Former, a multimodal module that bridges visual and textual representations at the hierarchy level for the different tasks (detection, classification, segmentation, morphology, mask language modeling and visual question answer) at different granularity. Extensive experiments demonstrate that Uni-Hema achieves comparable or superior performance to train on a single-task and single dataset models, across diverse hematological tasks, while providing interpretable, morphologically relevant insights at the single-cell level. Our framework establishes a new standard for multi-task and multi-modal digital hematopathology. The code will be made publicly available.
I Introduction
@@
  • We have designed Hema-Former, a multi-modal feature fusion mechanism. It inputs feature representations of text and image at different hierarchical levels, and using attention mechanism generates features of different granularity appropriate for different tasks. These different feature generation strategies are implemented through learnable 4 sub-modules. (Section III-B2)
  • We combine 46 diverse hematology datasets comprising 11 segmentation $(\approx 221K)$, 18 detection $(\approx 85K)$, and 17 classification $(\approx 350K)$ datasets, along with curated 22K Question Answer pairs and 7K masked language modeling samples, establishing the most extensive multi-modal corpus for digital Hematopathalogy to date.
@@ II Related Work
Vision Language Models (VLM’s): The emergence of VLM’s has bridged visual and textual modalities, enabling joint reasoning across computer vision and natural language processing tasks. Models such as CLIP [radford2021learning], ALIGN [jia2021scaling], and CoCa [yu2022coca] have pioneered task-agnostic pretraining by leveraging large-scale image–caption pair datasets to learn shared representation spaces, achieving strong generalization across diverse downstream tasks. Despite these advances, the lack of large-scale, high-quality image–text datasets specific to the sub-domains of pathology continues to constrain their generalizability and downstream applicability. Pathology oriented frameworks like CONCH [lu2023towards], Transpath [wang2021transpath], and Pathology foundation model (PFM) [ochi2024pathology] introduce domain adaptation strategies to improve multi-modal alignment for tissue slides. However, their transferability to hematology remains limited due to hematology-specific data scarcity, cells morphological diversity, and the fine-grained cellular variations that different from pathology imaging modalities. Foundation model: In the biomedical imaging domain, foundation models represent an effort to move beyond narrow task formulations by establishing transferable visual representations that can support a range of downstream analyses from tissue segmentation to cellular morphology classification. Approaches such as MedSAM [MedSAM], DinoBloom [koch2024dinobloom], and RedDino [zedda2025reddino] demonstrate the growing potential of large-scale pretraining in medical and hematology contexts. However, these models remain limited to single-cell analysis and textual integration, lacking unified multimodal reasoning across detection and contextual reasoning. While multimodal foundation models (e.g., UNI [chen2024towards], PFM [ochi2024pathology]) emphasize image–text retrieval but lake in fine-grained biomedical understanding. Unified model: Recent advancements have accelerated the development of unified models capable of performing multiple tasks within a single framework. Models such as DINO-X [ren2024dino], Uni-3DL [li2023uni3dl], and Uni-Perceiver v2 [li2023uni] exemplify this trend by incorporating multimodal inputs and modality-agnostic transformers to enable scalable, task-agnostic learning across diverse domains. However, in the medical imaging domain, unified architectures remain limited. Existing approaches, such as CelloType [pang2025cellotype] and LeukemiaAttri [rehman2024large] demonstrate progress toward multi-task learning by jointly performing dual tasks, but still rely on task-specific single dataset. Nonetheless, these methods remain constrained to specific diseases with annotated multitask datasets and lack the broader adaptability required for a unified model for multi-task, multi-disease, multimodal hematological analysis. To address the above mentioned limitation of VLMs, foundation and unfied models for hematopathology, our proposed unified framework leverages diverse datasets to enable integrated learning against multiple tasks and hematology diseases.
@@
III Methodology
III-A Dataset Formation
To train the unified model, we utilized 46 publicly available hematology datasets encompassing both single-cell and full-field microscopic images, covering a wide range of disease categories, including malaria, leukemia, anemia, sickle cell disease, and healthy samples. For segmentation or detection tasks, we generate a textual prompt for each image in the form: “This image is for the detection of <type of disease > of cells.” For Question Answer, the prompt starts with Q: and for Masked language modeling, the prompt starts with mask: tokens. The single-cell classification relied solely on visual features. The details of datasets is given in Section IV-A.
III-B Uni-Hema
@@
III-B1 Image and Text Encoders:
@@ III-B2 Hema-Former:
III-B3 Image and Text Decoders:
@@ IV-A Datasets
TABLE I: Performance of Uni-Hema across multiple tasks. Baseline methods are trained separately for each task and for each dataset. Uni-Hema, on the other hand, is trained only once and tested on the tasks without any re-training or fine-tuning. A “–” symbol indicates that a baseline model does not support the corresponding task. Values in bold and underlined denote the best and second-best results.
@@ IV-B Implementation details
Step 1: Pre-train the image backbone and encoder on single-cell (SC) classification datasets for 24 epochs with a batch size of 32. Step 2: Pre-train the text encoder and decoder on semi-synthetic masked language modeling (MLM) and question-answering (QA) datasets for the medical context learning. Step 3: Jointly train the vision modules of the Uni-Hema on classification, segmentation, and detection datasets using two images per task per batch (total batch size = 6) for 12 epochs, updating the visual modules, including the TGR, SCFE, and QGMF modules, simultaneously. The rest of the modules ($\mathbf{E}_{\mathbf{T}}$ , $\mathbf{D}_{\mathbf{T}}$, CMF) are freeze. Step 4: Fine-tune the image decoder for 12 epochs on all detection and microscopy-field-of-view (FoV) morphology datasets. Step 5: Fine-tune the query-guided mask former (QGMF) on the complete segmentation dataset for 8 epochs. Step 6: Train the cross-modal fusion (CMF) and $\mathbf{D}_{\mathbf{T}}$ on combined VQA and V-MLM datasets for 8 epochs, all other modules and submodules freeze. The complete training process requires approximately 8 days on a single NVIDIA RTX 4090 GPU (the best available resource that we have). The details of the training configuration for each step are provided in the Supplementary material.
IV-C Results and Discussion
@@
Detection Results: We compare the performance of cell detection across multiple datasets with previous state-of-the-art fully supervised methods as shown in Table IV-A. The results demonstrate that our proposed unified method achieves superior or comparable performance on cell detection tasks across various diseases, including leukemia, malaria, sickle-cell, and normal individuals. For sickle cell and parasite detection datasets, the number of training samples is very limited, making it difficult to learn dataset-specific patterns within a large corpus. In contrast, when training on individual datasets, the reduced complexity allows the model to learn more effectively.
@@
Experimental results Table IV-A demonstrate that Uni-Hema has superior performance compared to existing SOTA methods on both 21-class BMC  [matek2021highly], and 5-class Rabin  [kouzehkanan2021raabin] dataset.
Segmentation Results: To comprehensively assess segmentation performance, we compare Uni-Hema with leading fully supervised methods trained and tested on the particular dataset, including U-Net [ronneberger2015u], TransNetR [jha2024transnetr], and NanoNet [jha2021nanonet], as shown in Table IV-A. Our approach achieves competitive results, as reflected by the Dice score. Notably, the reported performance corresponds to a single unified model evaluated across all datasets, without any task-specific retraining.
Instead of a SOTA decoder for upsampling, we rely on simple mask upscaling during training (step 3) and a small upsampler during fine-tuning (step 5) due to resource constraints. Although this may slightly affect edge sharpness, our results remain comparable to the state-of-the-art TransNetR method. A qualitative comparison with TransNetR [jha2024transnetr] is illustrated in Figure 4.
Morphology Results: Cellular morphology across the microscopy field of view (FoV) is evaluated using the dataset from [rehman2024large] which contains 6 morphological attributes. As shown in Table IV-A, our method demonstrates superior performance compared to the corresponding SOTA multi-task, AttriDet [rehman2024large] method.
TABLE II: Comparison on unseen datasets: Uni-Hema demonstrates cross-domain adaptability across multiple tasks.
@@
Malaria [loddo2018mp] [shams2024bio] - - - - 78.5
@@
BBBC041Seg [tsutsui2023wbcatt] - - - - 86.2
@@
QA and MLM Results: To broaden the capability of our proposed unified model, we incorporate two text-aligned visual understanding tasks, single cell visual Question Answering (VQA) and full field-of-view Visual Masked Language Modeling (VMLM), into the training framework. As shown in Table IV-A, our UNI-Hema demonstrates strong performance across both tasks, achieving BLEU-4 scores of 56.4 for VQA, and 79.8 for V-MLM, respectively. These results highlight the model’s ability to bridge textual reasoning, including the visual interpretation in hematological contexts. Qualitative examples in Figure 4 further illustrate the model’s predictions.
Results on unseen datasets: We evaluate Uni-Hema on unseen datasets, and the results are shown in Table IV-C. For the classification task, we follow DinoBloom [koch2024dinobloom] experimental setup. Uni-Hema demonstrates overall superior performance withmean F1-score 90.8 . For cell detection task, it achieves a reasonable mAP@50 of 54.7 on the Bio-Net dataset [shams2024bio] and 78.5 on the Malaria dataset [loddo2018mp]. Furthermore, for segmentation on unseen datasets, Uni-Hema attains a Dice score of 86.2, highlighting its strong robustness across multiple tasks. For single-cell morphological analysis, we follow the WBCAtt pipeline [tsutsui2023wbcatt], treating the dataset as unseen for all feature extractor models [oquab2023dinov2, simeoni2025dinov3, koch2024dinobloom, he2016deep] listed in Table IV-C. The results show that Uni-Hema achieves state-of-the-art performance on single-cell morphology prediction.

### v2

@@ Abstract
Digital hematopathology requires cell-level analysis across diverse disease categories, including malignant disorders (e.g., leukemia), infectious conditions (e.g., malaria), and non-malignant red blood cell disorders (e.g., sickle cell disease). Whether single-task, vision-language, WSI-optimized, or single-cell hematology models, these approaches share a key limitation: they cannot provide unified, multi-task, multi-modal reasoning across the complexities of digital hematopathology. To overcome these limitations, we propose Uni-Hema, a multi-task, unified model for digital hematopathology integrating detection, classification, segmentation, morphology prediction, and reasoning across multiple diseases. Uni-Hema leverages 46 publicly available datasets, encompassing over 700K images and 21K question-answer pairs, and is built upon Hema-Former, a multimodal module that bridges visual and linguistic representations at the hierarchy level for the different tasks (detection, classification, segmentation, morphology, mask language modeling and visual question answer) at different granularity. Extensive experiments demonstrate that Uni-Hema achieves comparable or superior performance to train on a single-task and single dataset models, across diverse hematological tasks, while providing interpretable, morphologically relevant insights at the single-cell level. Our framework establishes a new standard for multi-task and multi-modal digital hematopathology. The code will be made publicly available.
1 Introduction
@@
  • We have designed Hema-Former, a multi-modal feature fusion mechanism. It inputs feature representations of text and image at different hierarchical levels, and using attention mechanism generates features of different granularity appropriate for different tasks. These different feature generation strategies are implemented through learnable 4 sub-modules. (Section 3.2.2)
  • We combine 46 diverse hematology datasets comprising 11 segmentation $(\approx 221K)$, 18 detection $(\approx 85K)$, and 17 classification $(\approx 320K)$ datasets, along with curated 22K Question Answer pairs and 7K masked language modeling samples, establishing the most extensive multi-modal corpus for digital Hematopathalogy to date.
@@ 2 Related Work
Vision Language Models (VLM’s): The emergence of VLM’s has bridged visual and textual modalities, enabling joint reasoning across computer vision and natural language processing tasks. Models such as CLIP [radford2021learning], ALIGN [jia2021scaling], and CoCa [yu2022coca] have pioneered task-agnostic pretraining by leveraging large-scale image–caption pair datasets to learn shared representation spaces, achieving strong generalization across diverse downstream tasks. Despite these advances, the lack of large-scale, high-quality image–text datasets specific to the sub-domains of pathology continues to constrain their generalizability and downstream applicability. Pathology oriented frameworks like CONCH [lu2023towards], Transpath [wang2021transpath], and Pathology foundation model (PFM) [ochi2024pathology] introduce domain adaptation strategies to improve multi-modal alignment for tissue slides. However, their transferability to hematology remains limited due to hematology-specific data scarcity, cells morphological diversity, and the fine-grained cellular variations that different from pathology imaging modalities.
Foundation model: In the biomedical imaging domain, foundation models represent an effort to move beyond narrow task formulations by establishing transferable visual representations that can support a range of downstream analyses from tissue segmentation to cellular morphology classification. Approaches such as MedSAM [MedSAM], DinoBloom [koch2024dinobloom], and RedDino [zedda2025reddino] demonstrate the growing potential of large-scale pretraining in medical and hematology contexts. However, these models remain limited to single-cell analysis and textual integration, lacking unified multimodal reasoning across detection and contextual reasoning. While multimodal foundation models (e.g., UNI [chen2024towards], PFM [ochi2024pathology]) emphasize image–text retrieval but lake in fine-grained biomedical understanding.
Unified model: Recent advancements have accelerated the development of unified models capable of performing multiple tasks within a single framework. Models such as DINO-X [ren2024dino], Uni-3DL [li2023uni3dl], and Uni-Perceiver v2 [li2023uni] exemplify this trend by incorporating multimodal inputs and modality-agnostic transformers to enable scalable, task-agnostic learning across diverse domains. However, in the medical imaging domain, unified architectures remain limited. Existing approaches, such as CelloType [pang2025cellotype] and LeukemiaAttri [rehman2024large] demonstrate progress toward multi-task learning by jointly performing dual tasks, but still rely on task-specific single dataset. Nonetheless, these methods remain constrained to specific diseases with annotated multitask datasets and lack the broader adaptability required for a unified model for multi-task, multi-disease, multimodal hematological analysis. To address the above mentioned limitation of VLMs, foundation and unfied models for hematopathology, our proposed unified framework leverages diverse datasets to enable integrated learning against multiple tasks and hematology diseases.
@@
3 Methodology
3.1 Dataset Formation
To train the unified model, we utilized 46 publicly available hematology datasets encompassing both single-cell and full-field microscopic images, covering a wide range of disease categories, including malaria, leukemia, anemia, sickle cell disease, and healthy samples. For segmentation or detection tasks, we generate a textual prompt for each image in the form: “This image is for the detection of <type of disease > of cells.” For Question Answer, the prompt starts with Q: and for Masked language modeling, the prompt starts with mask: tokens. The single-cell classification relied solely on visual features. The details of datasets is given in Section 4.1.
3.2 Uni-Hema
@@
3.2.1 Image and Text Encoders:
@@ 3.2.2 Hema-Former:
3.2.3 Image and Text Decoders:
@@ 4.1 Datasets
Table 1: Performance of Uni-Hema across multiple tasks. Baseline methods are trained separately for each task and for each dataset. Uni-Hema, on the other hand, is trained only once and tested on the tasks without any re-training or fine-tuning. A “–” symbol indicates that a baseline model does not support the corresponding task. Values in bold and underlined denote the best and second-best results.
@@ 4.2 Implementation details
Step 1: Pre-train the image backbone and encoder on single-cell (SC) classification datasets for 24 epochs with a batch size of 32. Step 2: Pre-train the text encoder and decoder on semi-synthetic masked language modeling (MLM) and question-answering (QA) datasets for the medical context learning. Step 3: Jointly train the vision modules of the Uni-Hema on classification, segmentation, and detection datasets using two images per task per batch (total batch size = 6) for 12 epochs, updating the visual modules, including the TGR, SCFE, and QGMF modules, simultaneously. The rest of the modules ($\mathbf{E}_{\mathbf{T}}$ , $\mathbf{D}_{\mathbf{T}}$, CMF) are freeze. Step 4: Fine-tune the image decoder for 12 epochs on all detection and microscopy-field-of-view (FoV) morphology datasets. Step 5: Fine-tune the query-guided mask former (QGMF) on the complete segmentation dataset for 24 epochs. Step 6: Train the cross-modal fusion (CMF) and $\mathbf{D}_{\mathbf{T}}$ on combined VQA and V-MLM datasets for 8 epochs, all other modules and submodules freeze. The complete training process requires approximately 8 days on a single NVIDIA RTX 4090 GPU (the best available resource that we have). The details of the training configuration for each step are provided in the Supplementary material.
4.3 Results and Discussion
@@
Detection Results: We compare the performance of cell detection across multiple datasets with previous state-of-the-art fully supervised methods as shown in Table 4.1. The results demonstrate that our proposed unified method achieves superior or comparable performance on cell detection tasks across various diseases, including leukemia, malaria, sickle-cell, and normal individuals. For sickle cell and parasite detection datasets, the number of training samples is very limited, making it difficult to learn dataset-specific patterns within a large corpus. In contrast, when training on individual datasets, the reduced complexity allows the model to learn more effectively.
@@
Experimental results Table 4.1 demonstrate that Uni-Hema has superior performance compared to existing SOTA methods on both 21-class BMC  [matek2021highly], and 5-class Rabin  [kouzehkanan2021raabin] dataset.
Segmentation Results: To comprehensively assess segmentation performance, we compare Uni-Hema with leading fully supervised methods trained and tested on the particular dataset, including U-Net [ronneberger2015u], TransNetR [jha2024transnetr], and NanoNet [jha2021nanonet], as shown in Table 4.1. Our approach achieves competitive results, as reflected by the Dice score. Notably, the reported performance corresponds to a single unified model evaluated across all datasets, without any task-specific retraining.
Instead of a SOTA decoder for upsampling, we rely on simple mask upscaling during training (step 3) and a small upsampler during fine-tuning (step 5) due to resource constraints. Although this may slightly affect edge sharpness, our results remain comparable to the state-of-the-art TransNetR method. A qualitative comparison with TransNetR [jha2024transnetr] is illustrated in Figure 4, while additional comparisons with other methods [ronneberger2015u, oktay2018attention, jha2021nanonet] are provided in the Supplementary Material.
Morphology Results: Cellular morphology across the microscopy field of view (FoV) is evaluated using the dataset from [rehman2024large] which contains 6 morphological attributes. As shown in Table 4.1, our method demonstrates superior performance compared to the corresponding SOTA multi-task, AttriDet [rehman2024large] method.
Table 2: Comparison on unseen datasets: Uni-Hema demonstrates cross-domain adaptability across multiple tasks.
@@
Malaria [loddo2018mp, shams2024bio] - - - - 78.5
@@
BBBC041Seg [tsutsui2023wbcatt, depto2021automatic] - - - - 86.2
@@
QA and MLM Results: To broaden the capability of our proposed unified model, we incorporate two text-aligned visual understanding tasks, single cell visual Question Answering (VQA) and full field-of-view Visual Masked Language Modeling (VMLM), into the training framework. As shown in Table 4.1, our UNI-Hema demonstrates strong performance across both tasks, achieving BLEU-4 scores of 56.4 for VQA, and 79.8 for V-MLM, respectively. These results highlight the model’s ability to bridge textual reasoning, including the visual interpretation in hematological contexts. Qualitative examples in Figure 4 further illustrate the model’s predictions.
Results on unseen datasets: We evaluate Uni-Hema on unseen datasets, and the results are shown in Table 4.3. For the classification task, we follow DinoBloom [koch2024dinobloom] experimental setup. Uni-Hema demonstrates overall superior performance withmean F1-score 90.8 . For cell detection task, it achieves a reasonable mAP@50 of 54.7 on the Bio-Net dataset [shams2024bio] and 78.5 on the Malaria dataset [loddo2018mp]. Furthermore, for segmentation on unseen datasets, Uni-Hema attains a Dice score of 86.2, highlighting its strong robustness across multiple tasks. For single-cell morphological analysis, we follow the WBCAtt pipeline [tsutsui2023wbcatt], treating the dataset as unseen for all feature extractor models [oquab2023dinov2, simeoni2025dinov3, koch2024dinobloom, he2016deep] listed in Table 4.3. The results show that Uni-Hema achieves state-of-the-art performance on single-cell morphology prediction.